OpenStudy (cderaalien):

I already know the solution, but how did the expression get simplified? n - 4 + 5n = 20

pooja195 (pooja195):

Start by combining like terms 5n+n=?

OpenStudy (cderaalien):

6n

pooja195 (pooja195):

Good 6n−4=20 add 4 to both sides 20+4=?

OpenStudy (cderaalien):

so, do I start by subtracting n from both sides?

OpenStudy (cderaalien):

* adding

pooja195 (pooja195):

Not quite we should always look for things we can combine it won't always be adding when dealing with negative numbers or subtraction signs you'll subtract

OpenStudy (cderaalien):

Oh!! We're adding like terms, I see now.

pooja195 (pooja195):

Yes! :-) Ok Good 6n−4=20 add 4 to both sides 20+4=?

OpenStudy (cderaalien):

24 then will we divide by 6 from both sides

OpenStudy (cderaalien):

on*

pooja195 (pooja195):

Yes perfect :)

OpenStudy (cderaalien):

which = 4 !! :) Thank you!! :)

OpenStudy (mathmale):

the -4 on the left and the 20 on the right are "like terms" in that they're both constants. Adding 4 to both sides of your equation results in 6n = What?

pooja195 (pooja195):

yw ^_^ and n=4

pooja195 (pooja195):

not just 4 :P

OpenStudy (cderaalien):

you guys are awesome

pooja195 (pooja195):

so are you :-) Welcome to openstudy!

OpenStudy (cderaalien):

Thank you lol.

pooja195 (pooja195):

yw ^_^